Logistics Solutions provides order fulfillment services for dot.com merchants. The company maintains warehouses that stock items carried by its dot.com clients. When a client receives an order from a customer, the order is forwarded to Logistics Solutions, which pulls the item from storage, packs it, and ships it to the customer. The company uses a predetermined variable overhead rate based on direct labor-hours. In the most recent month, 170,000 items were shipped to customers using 7,100 direct labor-hours. The company incurred a total of $23,430 in variable overhead costs. According to the company’s standards, 0.03 direct labor-hours are required to fulfill an order for one item and the variable overhead rate is $3.35 per direct labor-hour. Required: 1. What is the standard labor-hours allowed (SH) to ship 170,000 items to customers? 2. What is the standard variable overhead cost allowed (SH × SR) to ship 170,000 items to customers? 3. What is the variable overhead spending variance? 4. What is the variable overhead rate variance and the variable overhead efficiency variance?
Answer #1

Standard hours allowed = 170,000*0.03

= 5100

Standard cost allowed = 5100*3.35

= 17,085

Voh spending variance = 17,085 - 23,430

= 6345 U

Voh rate variance = (SR-AR) *AH

= (3.35 - 23,430/7100)*7100

= 355 F

Voh efficiency variance = (SH-AH)*SR

= (5100 - 7100)*3.35

= 6700 U